OpenStudy (hy123):

Find the critical numbers of the function. (Enter your answers as a comma-separated list. Use n to denote any arbitrary integer values. If an answer does not exist, enter DNE.) g(θ) = 8θ − 2 tan θ

OpenStudy (hy123):

@ganeshie8

OpenStudy (hy123):

Hi?

OpenStudy (anonymous):

critical points are the points where 1. function is not defined or 2. derivative of function is zero or 3. derivative of function is not defined

OpenStudy (hy123):

Yes, but im confused on how to work it.

OpenStudy (anonymous):

function is defined for all values. g'(θ) = 8 - 2 sec^2 θ 0 = 8 - 2 sec^2 θ 2 sec^2 θ = 8 sec^2 θ = 4 secθ = 2 or secθ = -2 => θ = n(pi) + 60degrees or θ = n(pi) - 60degress

OpenStudy (anonymous):

understood till there?

OpenStudy (hy123):

Yes

OpenStudy (anonymous):

So, g'(θ) becomes zero at those values...Now see where the g'(θ) is not defined. It is not defined when sec^2 θ is not defined And, sec^2 θ is not defined at θ = (2n+1)pi/2

OpenStudy (anonymous):

So, all those values of θ are called critical points

OpenStudy (anonymous):

got it?

OpenStudy (hy123):

I see that... So there it is undefined at both (2n+1)pi/2 and 2n-1)pi/2?

OpenStudy (anonymous):

take only (2n+1)pi/2 or (2n-1)pi/2..only one of them and put n = any integer...

OpenStudy (hy123):

So they are both similar?

OpenStudy (anonymous):

Yes..they are same for different values of n

OpenStudy (hy123):

Okay thank you.

OpenStudy (anonymous):

welcome:)